OpenStudy (anonymous):

For the given functions f and g, find the requested composite function value. f(x)= 2x+9; g(x)=(-1/x); find (g*f)(3) please help meee!

OpenStudy (anonymous):

\[f(x)=2x+9\] \[g(x)=-\frac{1}{x}\] \[g\circ f(x)=g(f(x))=g(2x+9)=-\frac{1}{2x+9}\]